OpenStudy (anonymous):

To four decimal places, what is the approximate solution of e^2x + 3 = 7? 1.386 0.6931 2.0000 2.4730

OpenStudy (campbell_st):

so is the equation 1. \[e^{2x} + 3 = 7\] or 2. \[e^{2x + 3} = 7\]

OpenStudy (anonymous):

1

OpenStudy (campbell_st):

ok... so subtract 3 from both sides... \[e^{2x} = 4\] next take the base e log of both sides \[\ln(e^{2x}) = \ln(4)\] apply the low law for powers \[2x \times \ln(e) = \ln(4)\] now you need to know that that the log of the base is 1 \[\log_{a}(a) = 1\] in this question you have \[\log_{e} (e) = \ln(e) = 1\] so you problem now is \[2x \times 1 = \ln(4)\] you should be able to solve it from here

OpenStudy (anonymous):

Okay ill try it

OpenStudy (campbell_st):

and check you answer by substituting back in...

OpenStudy (anonymous):

I cant figure it out. Can you show me farther?

OpenStudy (campbell_st):

ok... so you're happy \[2x = \ln(4)\] then divide both sides by 2 \[x = \frac{\ln(4)}{2}\] just evaluate that...